OpenStudy (anonymous):

if v1=(2 -5) and v2=(4 -3) then the angle between the two vectors is ?

OpenStudy (jdoe0001):

$$ cos(\theta)=\cfrac{\vec{v_1} \times \vec{v_2}}{||v_1|| \times ||v_2||}\\ ||v_1|| = \sqrt{x_1^2+y_1^2}\\ ||v_2|| = \sqrt{x_2^2+y_2^2} $$ once that's done, get the \(\large cos^{-1}\) of both, and that'd be your angle :)